OpenStudy (anonymous):

can you show me the steps for 5x+6=2x+x+2?

OpenStudy (anonymous):

5x+6=2x+x+2 5x+6=3x+2 carry 3x over the equal sine and 6 over the other side and change - to + and + to - so: 5x-3x=2-6 2x= -4 divide both sides by two x= -2